### v2

# Low Temperature Properties of Low-Loss Macroscopic Lithium Niobate Bulk Acoustic Wave Resonators

Abstract
We investigate gram scale macroscopic bulk acoustic wave (BAW) resonators manufactured from plates of piezoelectric lithium niobate. The intrinsic competing loss mechanisms were studied at cryogenic temperature through precision measurements of various BAW modes. Exceptional quality factors were measured for the longitudinal BAW modes in the 1-100 MHz range, with a maximum quality factor of 8.9 million, corresponding to a quality factor $\times$ frequency product of 3.8 ×$\times 10^{14}$ Hz. Through measurements of the acoustic response to a strong drive tone, anomalous self induced absorption and transparency effects are observed. We show that such observations can be explained by microscopic impurities and defect sites in the crystal bulk by the use of a non linear model of acoustic dissipation. The losses associated with these defects provide the ultimate limit of resonator performance, which could be improved in the future if more pure samples were available.
I Introduction
Bulk acoustic wave (BAW) resonators are crystalline devices engineered to sustain long-lived elastic vibrations propagating along the wave vector aligned with the crystal bulk. Historically, these devices have primarily been used for frequency standards and timing applications due to their inherent frequency stability and low acoustic loss [1]. Beyond engineering applications, the capability to support low-loss phonons makes them desirable for various tests of fundamental physics. For example, BAW phonons can be utilised for quantum information storage and computation in the field of quantum acoustic dynamics [2, 3, 4, 5], tests of quantum mechanical principles on large scales [6, 7], platforms for sensitive optomechanics [8] as well as sensors and detectors for beyond standard model physics, including searches for dark matter [9, 10], high frequency gravitational waves [11, 12, 13, 14, 15] and tests of Lorentz invariance [16, 17].
Low-loss BAW resonators are commonly manufactured from high purity crystalline quartz. This is partly owing to the fact that quartz displays significant improvements to acoustic loss when cooled to cryogenic temperatures[18, 19, 20, 21], making such devices ideal for low-energy precision sensing applications, such as tests of fundamental physics. As the requirements for such physics tests become more stringent however, it is natural to further the technological platform by considering alternative piezoelectric materials with differing properties.
Figure 1: a) Diagram of experimental set up. Cryostat environment was held under vacuum at temperatures of either 300 K or 4 K, and 10mK depending on the required measurement. b) Simulated mode distribution of the fifth order longitudinal mode. Red (blue) denotes the positive (negative) displacement in the $z$ direction
Piezoelectric lithium niobate for example provides the vibrating media for tunable micro-ring resonators [22, 23], optical frequency combs [24], micro electromechanical system (MEMS) based resonators [25], devices for optical to microwave transduction [26, 27, 28, 29, 30] and acoustic quantum state operations [31, 32, 33]. This material presents an attractive candidate as a macroscopic BAW resonator due to a stronger piezoelectric coupling coefficient, and smaller loss tangent than crystalline quartz at room temperature. However, material properties display a strong temperature dependence, and are thus drastically different when cooled. This motivates further investigation into the performance of macroscopic lithium niobate BAWs at cryogenic temperatures, for the purpose of next generation tests of fundamental physics. Additionally, the study of BAW modes at cryogenic temperatures allows one to investigate fundamental properties of the material.
In this work the loss mechanisms of lithium niobate BAW resonators are investigating by precise measurement of the quality factor ($Q$) of their acoustic modes at cryogenic temperatures. In the following sections II and III, the measurement set up is introduced and the frequencies of various fundamental bulk vibration modes are identified. Section IV then derives a theoretical model for the total $Q$ of a BAW resonator subject to accumulative loss mechanisms. In sections V and VI the measured $Q$ of various resonator modes are presented as functions of frequency and temperature respectively. Multiple competing loss mechanisms are identified based upon these observations. In section VII, we report on a strong drive amplitude dependence on the measured losses. Such self induced absorption and transparency effects are confirmed to be due to microscopic impurities by the use of an appropriate model of non-linear dissipation. Lastly, section VIII discusses the performance of these devices when further cooled to the extreme sub-Kelvin temperature of 10 mK.
II Measurement Set Up
The resonator in question consist of a 2 mm thick Z-cut lithium niobate disc, 30 mm in diameter. This geometry was chosen to emulate the typical standard sizes of commercially available macroscopic quartz resonators. One surface is machined with a plano-convex finish specified by a radius of curvature of 100 mm in order to minimise phonon losses via anchoring of the crystal at its edges [34]. The crystal plate is placed inside a copper cavity where it is clamped along its outer edge, with two electrodes placed $\approx$ 1 mm from the crystal surface. Each electrode can then be connected to co-axial lines that run out of the cryostat. The line that carries the excitation signal incident upon the resonator contains -20 dB of attenuation at the cryostat plate. This ensures that 300 K thermal noise on the drive signal is suppressed when operating at 4 K. An amplifier is also placed on the output line such that the returning signal is raised further above the instrument noise floor.
In this configuration, as shown in figure 1, acoustic resonances can be excited by an electrical drive signal applied to the electrodes. By connecting each co-axial line to ports of a vector network analyser; the acoustic mode frequencies $f_{r}$ are identified as peaks of maximum power transmission $S_{21}$. The input drive signal was varied across the frequency spectrum with 1 Hz resolution such that narrow resonances can be identified. The resulting spectra are then compared to a finite element model such that the mode structure of strongly coupled resonances can be identified.
Figure 2: Measurement of a)$f\cdot Q$ product and b) intrinsic $Q$ as a function of frequency for various longitudinal modes in the crystal resonators at $T=4$ K. The measured intrinsic quality factors are represented in blue, the black trace and grey bands denote the best fit and 95% confidence interval to the model of equation (3), the red trace shows the idealistic performance attainable without the presence of Rayleigh scattering. And the dashed black line gives the intrinsic limit to losses observed in similar quartz devices [35].
III Mode Structure
Two polarisation families of acoustic modes can be observed in the crystal due to the trigonal symmetry axis of lithium niobate. Longitudinally polarised resonances with bulk displacement in the crystal Z direction are denoted as A modes, while shear wave polarised modes with dominant displacement in the x-y plane are denoted B. In an ideal system only modes with odd overtone number $n$ can be coupled to external electrical signals as even $n$ generates a symmetric electric potential across the crystal. However, in the samples measured, even order modes can be observed with weak coupling. This is ascribed to crystal imperfections and asymmetries allowing for slightly asymmetric electric potentials in the vicinity of $n=\mathrm{even}$ resonances.
The frequencies of strongly coupled modes found by measurement were matched to the frequencies given by solutions to a finite element modelling as in figure 1 b), the results of which are presented in table 1. Close agreement between experimental data and model solutions is found. All odd overtones from $n=3$, to $n=15$, in both A and B polarisation are recorded in the table. For higher frequencies, the finite element resolution of the model needs to be further increased in order to guarantee accuracy. However from the results provided in table 1, a simple linear extrapolation $f_{i}/n_{i}=f_{i}/n_{i}$ can be utilised to identify any $n_{ij}$ overtone, of frequency $f_{ij}$, for modes of the same polarisation family.
$X_{n,m,p}$ $f_{r}^{\mathrm{(model)}}$ $f_{r}^{\mathrm{(exp)}}$ $X_{n,m,p}$ $f_{r}^{\mathrm{(model)}}$ $f_{r}^{\mathrm{(exp)}}$
B5,0,0 4.08 4.20 A7,0,0 11.63 11.78
A3,0,0 4.90 4.75 B15,0,0 12.22 12.24
B7,0,0 5.70 5.793 A9,0,0 14.94 14.98
B9,0,0 7.34 7.40 A11,0,0 18.279 18.26
A5,0,0 8.31 8.33 A13,0,0 21.60 21.48
B11,0,0 8.96 9.02 A15,0,0 24.925 24.8
B13,0,0 10.60 10.64 - - -
Table 1: Frequencies in MHz of the fundamental modes found via finite element modelling $f_{r}^{\mathrm{(model)}}$, compared to the experimentally found frequencies $f_{r}^{\mathrm{(exp)}}$.
IV Theoretical Model of Acoustic Loss
Figure 3: $Q$ as a function of temperature $T$ is measured for the B13,0,0 mode and the resulting scaling relationship between losses $1/Q$ and temperature is shown. Inset shows the scaling for the A19,0,0 mode when cooled to 10 mK. Temperature scaling characteristic of TLS limited loss is identified, however the presence of strong absorption peaks in close proximity are also observed.
The dominant loss mechanisms in BAW resonators can be understood by measurement of the intrinsic quality factor of the acoustic modes $Q_{i}$. This quantity is the inverse of the total phonon losses and as such can be parameterised as the inverse sum of the quality factor associated with each competing loss mechanism. For bulk waves, the contributing losses to $Q_{i}$ can be due to clamping of the vibrational structure to external supports, thermoelastic damping, phonon-phonon scattering, and phonon-defect scattering. Phonon-electron scattering loss can be dominant in semi-conductor systems or at microwave frequencies, however it is negligible here.
For acoustic resonators at 300 K, phonon-phonon scattering loss is well described by the macroscopic Akheiser regime [36], which predicts a constant $Q\cdot f$ product. For lithium niobate the theoretical limit to $Q\cdot f$ as predicted by the Akheiser model is $Q\cdot f\approx 4.74\times 10^{14}$, although the largest room temperature value observed in the devices measured in this work is $1.12\times 10^{14}$ corresponding to a maximum $Q$ of $3.28\times 10^{6}$ However, at lower temperatures the life time of thermal phonons $\tau_{\mathrm{th}}$ increases such that phonon attenuation process undergoes a phase change into the microscopic description of Landau and Rumer [37]. In this regime, the attenuation displays a weaker frequency dependence. As a result $Q\cdot f$ begins to scale as $Q\cdot f\propto f$, and $Q$ also greatly increases with decreasing temperature. It is precisely this change to the fundamental limit of intrinsic acoustic loss that gives rise to the numerous applications of high overtones BAW devices at low temperature. For lithium niobate at 4 K phonon modes are well described by the Landau-Rumer model for $f\gtrsim 1.4$ MHz, which equivalently gives $12\pi f\tau_{\mathrm{th}}>$ 1.
Thermoelastic damping occurs in longitudinal bulk vibrations as the elastic deformation of the travelling wave couples to the thermal expansion tensor associated with local temperature perturbations. Thus longitudinal modes can dissipate energy through the materials thermal conductivity. In the regime where the acoustic wave propagation can be considered quasi-isothermal, the quality factor associated with thermoelastic damping may be written as [18, 38]
$Q_{\mathrm{TED}}=\frac{2\pi f}{f^{*}}\frac{\rho C^{s}c_{33}}{\lambda_{33}^{2}T}~{}~{}~{}\lambda_{33}:=2c_{12}\alpha_{11}+c_{33}\alpha_{33},$ (1)
where $\rho$ is the material density, $T$ is the reference temperature, $C^{s}$ is the specific heat capacity at constant strain, $c_{ij}$ is the isothermal elastic tensor and $\alpha_{ij}$ the thermal expansion tensor. The turning point frequency $f^{*}$ denotes the transition point between the adiabatic and isothermal propagation regimes, it is given by $f^{*}=V_{3}^{2}/2\pi\chi_{3}$ where $\chi_{3}$ is the material diffusivity. For lithium niobate at $T=4$ K the thermal conductivity and specific heat have been measured, but the elastic constants and expansion coefficients are unknown. Approximating the elastic constants from their room temperature values gives $f^{*}\approx 13$ MHz, with $f>f^{*}$ defining the isothermal regime.
For a convex curved vibrating plate, phonons are effectively trapped and confined to the centre, suppressing phonon escape through the support structure and thus limiting the clamping losses. However for low order modes this attenuation is still significant, and has been previously approximated [34] by considering the ratio of vibrational energy in a finite curved plate compared to the total energy of an infinite plate:
$Q_{\mathrm{Clamping}}=\left(1-\mathrm{erf}\left(\eta\sqrt{n}\right)^{2}\right)^{-1}.$ (2)
Where $n$ is the mode overtone number in the thickness direction and $\eta$ is a dimensionless trapping parameter [39] that should be greater than unity for well trapped phonons. It is clear from this equation that the clamping loss rapidly decreased with increasing $n$.
Acoustic attenuation due to either structural defects or microscopic impurities can be harder to quantify exactly, however scaling relationships can be predicted in some cases. Considering the standard tunnelling model of two level systems in amorphous glasses [40, 41], it can be shown [42] that the acoustic absorption by two level systems, potentially arising from non-crystalline impurities, results in a temperature dependent scaling of the acoustic losses $1/Q\propto T^{1/3}$. At lower temperatures populations of these two level systems can become unsaturated and have a dominant impact on total acoustic dissipation.
In addition to temperature scaling, Rayleigh-type scattering loss of phonons in metals and glasses is predicted to scale as $1/Q_{\mathrm{Rayleigh}}\propto f^{-3}$ [43]. This relationship however, has also been observed for high overtones in similar quartz piezoelectric BAW systems [35], suggesting the presence of small populations of impurity ions that cause amorphous behaviour at high frequencies.
Examining the measured intrinsic $Q$’s for various longitudinal overtones in lithium niobate presented in figure 2, the presence of multiple competing loss regimes can be identified. An appropriate model for the total losses in the system, as a function of frequency, can be written as
$\frac{1}{Q_{i}}=\frac{1}{Q_{\mathrm{Clamping}}}+\frac{1}{Q_{\mathrm{TED}}}+\frac{1}{Q_{\mathrm{Rayleigh}}}.$ (3)
V Frequency Dependence of Acoustic Loss
The intrinsic quality factor of bulk lithium niobate modes can be measured by fitting a Lorentzian model to the $S_{21}$ spectrum of a given mode and extracting its linewidth $\gamma$, such that $Q_{i}=f_{r}/\gamma$. The measurement uncertainty following this procedure is then quantified as the standard error associated to the fit parameters. As the motional resistance of these modes are far greater than the source impedance of the electrical measurement network, the modes are severely under coupled, and thus the measured $Q$ corresponds exactly to the intrinsic quality factor of the resonator $Q_{i}$. The measured values from this procedure are shown in figure 2. The cryogenic quality factors show consistent improvement from room temperature values across all frequencies, with a maximum recorded quality factor of $8.9\times 10^{6}$.
By substituting equations (1) and (2) and the functional form $Q_{\mathrm{Rayleigh}}=10^{a}f^{-3}$ into equation (3), a fit can be performed to the measured quality factors as a function of frequency with $\lambda_{33}$, $\eta$ and $a$ as fit parameters. The resulting best fit and confidence interval band can also be observed in figure 2.
The fit extracts a trapping of $\eta=2.2058\pm 0.24$ and $\lambda_{33}=2228.27\pm 73.67$ Pa K-1. Assuming for a moment isotopic thermal expansion such that $\alpha_{12}=\alpha_{33}$, an expansion coefficient of $5.63\times 10^{-8}~{}\mathrm{K}^{-1}$ would be attained. Comparing this value to the known coefficients at room temperature of $\alpha_{12}\approx 15\times 10^{-5}~{}\mathrm{K}^{-1}$ and $\alpha_{33}\approx 4\times 10^{-5}~{}\mathrm{K}^{-1}$ we see that the reduction of temperature reduces thermal expansion by roughly three orders of magnitude.Whilst the exact quantitative nature of this argument relies on a linear extrapolation of the known temperature variation in elastic tensor components at $T=273$ K [44], as well as the fact that the there is no reason to believe that thermal expansion is isotropic; the determined vast reduction to the thermal expansion tensor is absolute. Indeed reductions to $\alpha_{ij}$ of similar order have been measured previously in piezoelectric quartz when cooled down to temperatures of 1 K [45].
The primary significance of this model is the accuracy in the predicted relationships between $Q$ and $f$ across three different regimes defined by competing dominant loss mechanisms. It is observed that thermoelastic damping limits device performance at cryogenic temperatures for some frequencies of interest, a characteristic that is not observed in quartz [18] due to the smaller diffusivity and phonon velocities when compared to the measured values for lithium niobate at low temperature [46]. Additionally, at high frequencies performance becomes limited by Rayleigh scattering of heavy impurity ions. This impurity loss represents a non-fundamental engineering barrier that could be overcome by better crystal purity, thus an additional idealistic loss model has also been plotted on figure 2 that shows the theoretical device performance of a pure crystal. Even assuming perfectly pure crystals, the intrinsic loss limit at 4 K in lithium niobate is still orders of magnitude below that of quartz resonators at the same temperature and in the presented frequency range. This is due to the stronger thermoelastic damping in lithium niobate when compared to quartz, which has been observed to follow the theoretical Landau-Rumer dissipation at this temperature and frequency range [35]. This observed limit for quartz is shown by the dashed black trace in figure 2 for direct comparison.
Typical impurity concentrations from the crystal growth procedure record a maximum impurity concentration of fluoride at $<50$ parts per million, with typical concentrations of most other elements at $\lesssim 1$ parts per million, suggesting that even very small populations of impurities greatly inhibit performance benefits at high frequencies. In order to mitigate Rayleigh scattering loss at these frequencies crystal samples would need ot be produced with impurity populations at the level of parts per billion, as has been seen in nano-mechanical devices [47].
VI Temperature Dependence of Acoustic Loss
The measurement of the quality factor’s temperature dependence, $Q(T)$ offers further insights into the dominant loss mechanisms, such as regimes where losses are primarily influenced by two-level systems (TLS) associated with impurity ions. By observing the scaling relationship of losses ($1/Q$) with temperature as shown for the B13,0,0 10.6 MHz mode on the main axes of figure 3, we see the presence of absorption peaks for defined temperature $\approx 28$ K and above. Such increased absorption in narrow temperature regions is due to the system entering an energy regime in which the thermal bath become resonant with phonons. These features have been observed in acoustic systems at low temperature previously [48], and are not indicative of TLS induced dissipation.
The characteristic scaling law expected for intrinsic losses due to the coupling of phonons to impurity TLS is theorised to follow $n=1/3$ [49, 18], however many experimental studies on low temperature acoustic systems [18, 50, 51, 52] have recorded a scaling slightly skewed from this exact value due to competing losses from multiple TLS ensembles and excess loss from other mechanisms such as resonant absorption. The regime between 10 and 20 K shows a temperature dependence of $1/Q\propto T^{0.34}$ that is reminiscent of TLS scattering. However, due to the presence of resonant absorption peaks in close proximity, and the loss saturation observed for $T<10$ K, it is hard to say with certainty if this effect pure TLS absorption, or multiple other competing mechanisms.
In the inset of figure 3, we show the further cooling of the $A_{19,0,0}$ 31 MHz mode down to $T=$10 mK, where a weaker temperature dependence was observed in the 1-4 K regime, most likely due to the presence of resonant absorption. At even lower temperatures the losses saturate and no further improvement is found, showing that the dominating dissipation channel cannot be ascribed to phonon-phonon scattering.
VII Self Induced Transparency and Absorption
Figure 4: Measured $S_{21}$ transmission spectra is shown for the $19^{\mathrm{th}}$ longitudinal overtone mode as a function of drive power. The drive power is normalised to the critical power value $P_{c}$, at which point the crystal non-linear regime becomes dominant and the mode begins to demonstrate self induced absorption.
Figure 5: Numerical solutions to equation (4) determined via harmonic balance methods are presented. a) The frequency response spectrum of a singular solution branch is plotted as a function of increasing drive force. The drive is normalised by the threshold level that induces absorption $F_{0}$. b) The parameter space in which stable solutions $n$ exist, is shown. Each of the three stable solution families $n$ are distinguished by a corresponding colour for a given frequency $f$ and force $F$. These simulations demonstrate clear resemblance to the data of figure 4, in which the bi-stable region is symmetrical about resonance, and gives a drive dependent regime of increased absorption.
Investigating the lithium niobate resonators at large drive amplitudes, strongly non-linear regimes are observed. Utilising the same set up as described in section II, the resonator modes transmission spectrum were measured as a function of increasing drive power $P$. An example is presented In figure 4, where a non linear regime is reached for some critical power level $P_{c}$ in the $19^{\mathrm{th}}$ longitudinal overtone mode. After which the transmission collapses, and the frequency response enters a regime of increased dissipation, a phenomena known as self induced absorption. At even larger drive amplitudes, the response ultimately recovers during the inverse process of self induced transparency.
Bulk acoustic wave systems in general are known to enter non-linear regimes in which higher order terms in the vibrational amplitude $u$ begin to dominate system behaviour. A famous example of this is the existence of the Duffing non-linearity present in many acoustic systems [53, 54, 55]. This effect is characterised by an asymmetrical, amplitude dependent perturbation to the resonant frequency due to high order elastic constants generating terms cubic in $u$. However, the self induced absorption and transparency phenomena of figure 4 cannot be explained by a duffing model, as the discontinuities in frequency response occur symmetrically about resonance.
In so called ‘un-swept’ piezoelectric quartz BAW resonators that are yet to undergo a sweeping impurity removal process, various anomalous nonlinear effects are also observed that cannot be explained by the Duffing model [35]. The quartz devices exhibit discontinuities in their frequency response for certain drive amplitudes, van-der-pol like quadrature oscillations, as well as regimes of self induced transparency and absorption. These effects are reported to be due to the existence of heavy impurity ions in the crystalline bulk, prevalent to the levels of parts per million.
Self induced transparency and absorption has also previously been observed in acoustic media with artificially introduced defects [56, 57]. Here the defects were modelled as localised non linear dissipation terms introduced to the resonator’s equations of motion. These terms then introduce perturbations on top of the linear losses in certain regimes. In previous work with un-swept quartz analytical arguments were made to show that such a model could qualitatively reproduces the observed phenomena.
Taking the simplified dissipation model for the case of a singular defect[35], and considering a polynomial non-linear dissipation function $f(\dot{u}))$ [56], we can write the equation of motion
$\ddot{u}+2\gamma(1+f(\dot{u}))\dot{u}+\omega_{r}^{2}u=F\mathrm{cos}(\omega t),$ (4)
where $\omega$ is angular frequency, $\gamma$ parameterises linear dissipation, $F$ is the external driving force to the system, and $f(\dot{u}))$ is given by $f(\alpha)=-\beta_{1}\dot{\alpha}^{2}+\beta_{2}\dot{\alpha}^{4}$ with appropriate choices for the constants $\beta_{1,2}$.
This non linear model can then be solved utilising numerical harmonic balance methods in order to reproduce the anomalous self induced absorption effect. As the system is strongly-nonlinear, a variety of solution branches for $u$ are attained. In figure 5 a) the spectrum of normalised displacement amplitudes $u(f)$ is plotted for one of the solution branches, as a function of drive power $F$. The same symmetrical discontinuity about resonance can be observed as in figure 4, as the response collapses under self induced absorption due to the presence of non-linear dissipation that dominates above a threshold drive force. By visualising the parameter space for stable solutions to equation (4), plotted in figure 5 b), it is clear that the non-linear dissipation term causes switching between bi-stable states on both sides of the resonant peak. This recreates the characteristic self induced absorption phenomena. The accuracy of this model in reproducing the non-linear behaviour observed in figure 4 suggests that populations of heavy ions, to the level in parts per million, limit the performance of lithium niobate BAW resonators, as in un-swept quartz.
Figure 6: Measured quality factors at $T=4$ K from figure 2 are plotted with the quality factors for the same modes measured at $T=10$ mK as shown in red. Increased absorption is clearly observed in the sub-Kelvin regime.
VIII Performance at Sub-Kelvin Temperature
The cryostat system as described in sectionII can be further cooled from 4 K down to sub-Kelvin temperatures due to helium dilution. This provides an opportunity to explore the lithium niobate system at extreme cryogenic temperatures approaching the standard quantum limit. Taking a look at figure 6 the quality factors greatly suffer when cooled further to 10 mK. This is due to the increased presence of non linear dissipation channels, such as regimes dominated by self induced absorption. The stronger presence of non linear dissipation is provided by the thermal un-saturation of two level system states due to microsopic impurities. The presence of these states make it difficult to operate the system below nonlinear amplitude thresholds, resulting in general increases to dissipation.
IX Conclusion
Z-cut lithium niobate provides a promising material for macroscopic BAW resonators designed with the intention of testing areas of fundamental physics, demonstrating large quality factors for bulk acoustic modes. However, the existence of microscopic impurities sites in the crystalline bulk puts limitations on the ultimate performance for high overtones. Additionally, the strong thermoelastic damping of the material results in further attenuation at low frequencies in comparison to crystalline quartz. We demonstrate that the losses in our system are well understood, with intrinsic quality factors adhering to models predicted by theory. Future improvements to lithium niobate resonators can be made by reducing atomic impurity levels. Treatments such as DC sweeping, and improvements to crystal growth procedures have been successfully to produce higher quality material in other systems. Pending such engineering improvements lithium niobate BAW resonators will see high overtone quality factors tend towards the billions, reaching the intrinsic limit set by thermoelastic damping and phonon-phonon scattering. Such high quality factors will be beneficial to various fundamental physics experiments hoping to exploit the large mass macroscopic phonon modes provided by this system.